# Limitations of Data-Driven Spectral Reconstruction – Optics-Aware Analysis and Mitigation

### v2

@@ 5.2. Training with metamers
Figure 5. Training MST++ with metamers. It fails to combat fixed metamers and on-the-fly metamers, in particular on the spectral accuracy SAM.
As an example, we train MST++ and evaluate its validation performance over the training process. In Fig. 5, we show that it is no longer a good choice to use MRAE as the loss function (Cai et al., 2022b) and evaluation metric (Arad et al., 2022), because it is completely overwhelmed by metamers. Instead, we find that L1 loss is a more stable loss function, so we train the network with L1 loss for three cases, no metamer (easy), fixed metamer (medium), and on-the-fly metamer (difficult). Nevertheless, we can see in Fig. 5 that the network fails in particular for the spectral accuracy SAM.
Table 6. Performance comparison for training various networks with fixed and on-the-fly metamers.
@@
We also train all other candidate networks with fixed and on-the-fly metamers. The results are summarized in Table 6. Again, the same performance drop applies to all networks. Finally, we show the results of the top-performing network, MST++ on the CAVE, ICVL, and KAUST datasets in Table 7. (See Supplemental Material for more results). As before, the performance drops similarly in the presence of metamers.
5.3. Discussion
@@
The limitations of spectral estimation from RGB data are ultimately not overly surprising – after all the projection from the high dimensional spectral space to RGB invariably destroys scene information that can be difficult to recover. Spatial context from underrepresented data does not contribute to the spectral estimation, because such information remains the same for metamers. However, our experiments show that this is indeed an issue faced by the state-of-the-art methods, which so far went unnoticed due to the under-representation of metamers in the datasets. This shortcoming will also affect other uses of the same datasets, for example in the training of reconstruction methods for spectral computational cameras (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011). Metameric adversary helps to identify this overlooked issue, whereas metameric augmentation has not yet found its way for contribution. We will explore it further in the next section.
Table 7. Training with metamers for MST++ on the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets.
@@
6. Finding 3: The Aberration Advantage
As shown so far, the existing methods have difficulties distinguishing metamers in the ideal noiseless and aberration-free condition. In this section, we analyze what effect (if any) optical aberrations have on this situation, i.e., aberration-aware training (Yang et al., 2023). To this end we train the networks in a realistic imaging condition with moderate noise level (npe = 1000), lossless PNG format, and aberrations from the same double Gauss lens as before (Ichimura, 2021). In short, we simulate, through spectral ray tracing, the effect that an imperfect (i.e., aberrated) optical system has on the RGB image measured when observing a specific spectral scene. The details of this simulation can be found in the Supplemental Material.
In Fig. 6, we show an example with MST++ for the validation on SAM in two situations, one with fixed metamers, and the other with on-the-fly metamers. In each experiment, we compare the SAM difference with and without aberrations. As a reference, we also show the standard validation without metamers. As we can see, the realistic optical aberrations of the lens actually improve the spectral estimation in the presence of metamers as long as the aberrations are modeled in the training. With chromatic aberrations, the network can already distinguish fixed metamer pairs, achieving similar accuracy as the standard case. In the more aggressive case of on-the-fly metamers, chromatic aberrations also improve the spectral accuracy, compared with its no-aberration counterpart. Again, this aberration advantage holds for all datasets (Table 8). See Supplemental Material for details.
@@ 6. Finding 3: The Aberration Advantage
Figure 6. Chromatic aberrations improve spectral accuracy. Left: fixed metamers. Right: on-the-fly metamers.
Table 8. SAM metrics for MST++ on CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021).
@@
6.1. Discussion
To understand why optical aberrations help improve the reconstruction, consider the simulated images in Fig. 7. The left and middle images are simulations of RGB images for metameric scene pairs, with the difference image on the right. The different spectra of the two scenes are affected differently by the optical aberrations, and therefore, although the scenes are metamers of each other, the RGB images are in fact different. In effect, the optical aberrations have encoded spectral information into the RGB image, which the networks can learn to distinguish, lending credibility to PSF engineering methods for hyperspectral encoding (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011).
@@ 6.1. Discussion
Figure 7. Chromatic aberration induced informative color differences (right) as spectral cues for metamer pairs (left and middle).
There are, however, also detrimental consequences of this effect, namely (1) “baked-in” aberrations in spectral datasets and (2) limited generalization to arbitrary RGB image sources.
Regarding (1), it is worth noting that even expensive spectral cameras typically do not have diffraction limited optics, i.e., their optical systems are aberrated. In order to avoid overfitting to these aberrations, future datasets should ideally be collected with a large variety of different spectral cameras.
With respect to (2) we note that different RGB cameras produce very different aberrations, and often the aberrations also vary significantly across the image plane. Without further improvements on the reconstruction methods, this will result in a large device dependency and poor generalization of the methods.
7. Conclusion
@@
In particular the last point lends additional credibility to computational camera approaches for spectral imaging, including both PSF engineering approaches (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011) and compressed sensing methods like CASSI (Wagadarikar et al., 2008; Choi et al., 2017b). However, learned reconstruction methods for these approaches will suffer from the same dataset issues as the methods analyzed in this paper, making the collection of large-scale, diverse spectral image data a matter of urgency.
In order to realize the dream of spectral estimation from arbitrary RGB sources it is necessary to diversify the spectral contents in hyperspectral image datasets and embrace versatile spectral data augmentation methods to fully enable the power of networks in adaptation to whole families of aberrations.
Acknowledgements.
This work was supported by the King Abdullah University of Science and Technology Individual Baseline Funding. Seung-Hwan Baek was supported by Korea NRF (RS-2023-00211658) and Samsung Electronics.
@@ Acknowledgements.
Appendix A Image Formation Model
@@
(10) ${\mathbf{X}}=\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right],$
@@
(11) ${\mathbf{Q}}=\left[{\mathbf{q}}_{1},{\mathbf{q}}_{2},{\mathbf{q}}_{3}\right]=\left[\begin{matrix}q_{11}&q_{21}&q_{31}\\ q_{21}&q_{22}&q_{32}\\ \vdots&\ddots&\vdots\\ q_{K1}&q_{K2}&q_{K3}\\ \end{matrix}\right],$
@@
(12) ${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$
@@
(13) ${\mathbf{Y}}=\left[{\mathbf{y}}_{1},{\mathbf{y}}_{2},{\mathbf{y}}_{3}\right],$
@@
(14) ${\mathbf{w}}_{k}={\mathbf{A}}_{k}{\mathbf{x}}_{k},\quad k\in[1,2,\dots,K],$
@@
(15) ${\mathbf{W}}$ $=\left[{\mathbf{w}}_{1},{\mathbf{w}}_{2},\dots,{\mathbf{w}}_{K}\right]$
@@
(16) ${\mathbf{A}}=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right],$
@@
(17) ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right),$
@@
(18) ${\mathbf{A}}{\mathbf{X}}$ $=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right]\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right]$
@@
(19) ${\mathbf{Z}}={\mathbf{W}}{\mathbf{Q}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}}.$
@@
Appendix B Effect of clipping to non-negative values.

### v3

@@ V-B Training with metamers
Figure 5: Training MST++ with metamers. It fails to combat fixed metamers and on-the-fly metamers, in particular on the spectral accuracy SAM.
As an example, we train MST++ and evaluate its validation performance over the training process. In Fig. 5, we show that it is no longer a good choice to use MRAE as the loss function [14] and evaluation metric [7], because it is completely overwhelmed by metamers. Instead, we find that L1 loss is a more stable loss function, so we train the network with L1 loss for three cases, no metamer (easy), fixed metamer (medium), and on-the-fly metamer (difficult). Nevertheless, we can see in Fig. 5 that the network fails in particular for the spectral accuracy SAM.
TABLE VI: Performance comparison for training various networks with fixed and on-the-fly metamers.
@@
We also train all other candidate networks with fixed and on-the-fly metamers. The results are summarized in Table VI. Again, the same performance drop applies to all networks. Finally, we show the results of the top-performing network, MST++ on the CAVE, ICVL, and KAUST datasets in Table VII. (See Supplemental Material for more results). As before, the performance drops similarly in the presence of metamers.
V-C Discussion
@@
The limitations of spectral estimation from RGB data are ultimately not overly surprising – after all the projection from the high dimensional spectral space to RGB invariably destroys scene information that can be difficult to recover. Spatial context from underrepresented data does not contribute to the spectral estimation, because such information remains the same for metamers. However, our experiments show that this is indeed an issue faced by the state-of-the-art methods, which so far went unnoticed due to the under-representation of metamers in the datasets. This shortcoming will also affect other uses of the same datasets, for example in the training of reconstruction methods for spectral computational cameras [36, 9, 15]. Metameric adversary helps to identify this overlooked issue, whereas metameric augmentation has not yet found its way for contribution. It also underscores that, without side-channel information, no intrinsic property exists in the solely use of RGB images to distinguish between metamers, even when they are augmented. Note that this does not downplay the effects of metameric augmentation, since the problem formulation of reconstructing spectral information from RGB images in Eq. (1) is fundamentally limited. Once the problem is formulated in Eq. (2), metameric augmentation contributes to improving the network robustness. We will explore it further in the next section.
TABLE VII: Training with metamers for MST++ on the CAVE [67], ICVL [4], and KAUST [42] datasets.
@@
VI Finding 3: The Aberration Advantage
As shown so far, the existing methods have difficulties distinguishing metamers in the ideal noiseless and aberration-free condition. In this section, we analyze what effect (if any) optical aberrations have on this situation, i.e., aberration-aware training [66]. To this end we train the networks in a realistic imaging condition with moderate noise level (npe = 1000), lossless PNG format, and aberrations from the same double Gauss lens as before [33]. In short, we simulate, through spectral ray tracing, the effect that an imperfect (i.e., aberrated) optical system has on the RGB image measured when observing a specific spectral scene. The details of this simulation can be found in the Supplemental Material.
In Fig. 6, we show an example with MST++ for the validation on SAM in two situations, one with fixed metamers, and the other with on-the-fly metamers. In each experiment, we compare the SAM difference with and without aberrations. As a reference, we also show the standard validation without metamers. As we can see, the realistic optical aberrations of the lens actually improve the spectral estimation in the presence of metamers as long as the aberrations are modeled in the training. With chromatic aberrations, the network can already distinguish fixed metamer pairs, achieving similar accuracy as the standard case. In the more aggressive case of on-the-fly metamers, chromatic aberrations also improve the spectral accuracy, compared with its no-aberration counterpart. To ablate on the effects of aberrations and metameric augmentation, we also plot the result with aberrations, but without metameric augmentation (purple lines). It implies that with aberrations only, spectral accuracy cannot be improved, because it simply means to reconstruct the hyperspectral images with more blurred RGB images. Again, this aberration advantage holds for all datasets (Table VIII). See Supplemental Material for details.
@@ VI Finding 3: The Aberration Advantage
Figure 6: Chromatic aberrations improve spectral accuracy. Left: fixed metamers. Right: on-the-fly metamers.
TABLE VIII: SAM metrics for MST++ on CAVE [67], ICVL [4], and KAUST [42].
@@
VI-A Discussion
To understand why optical aberrations help improve the reconstruction, consider the simulated images in Fig. 7. The left and middle images are simulations of RGB images for metameric scene pairs, with the difference image on the right. The different spectra of the two scenes are affected differently by the optical aberrations, and therefore, although the scenes are metamers of each other, the RGB images are in fact different. In effect, the optical aberrations have encoded spectral information into the RGB image, which the networks can learn to distinguish, lending credibility to PSF engineering methods for hyperspectral encoding [36, 9, 15].
@@ VI-A Discussion
Figure 7: Chromatic aberration induced informative color differences (right) as spectral cues for metamer pairs (left and middle).
TABLE IX: Generalization analysis on the effects of the proposed metameric augmentation for various spectral encoding schemes.
Trained on Validated on No metameric augmentation With metameric augmentation
RGB2HS CASSI RGB2HS RGB2HS + aberrations CASSI
PSNR SAM PSNR SAM PSNR SAM PSNR SAM PSNR SAM
ARAD1K CAVE std 29.9 0.359 30.6 0.270 22.4 0.692 31.1 0.363 31.7 0.257
CAVE met 27.3 0.510 24.2 0.382 29.9 0.296 37.9 0.160 27.4 0.295
ICVL std 24.5 0.090 39.7 0.048 35.5 0.168 34.4 0.164 39.9 0.046
ICVL met 24.5 0.497 27.8 0.314 33.9 0.290 34.1 0.311 32.8 0.195
KAUST std 23.5 0.512 41.9 0.099 32.7 0.253 35.5 0.247 44.3 0.078
KAUST met 25.3 0.775 32.8 0.284 34.0 0.220 39.9 0.073 37.2 0.184
There are, however, also detrimental consequences of this effect, namely (1) “baked-in” aberrations in spectral datasets and (2) limited generalization to arbitrary RGB image sources. Regarding (1), it is worth noting that even expensive spectral cameras typically do not have diffraction-limited optics, i.e., their optical systems are aberrated. In order to avoid overfitting to these aberrations, future datasets should ideally be collected with a large variety of different spectral cameras. Regarding (2) we note that different RGB cameras produce very different aberrations, and often the aberrations also vary significantly across the image plane. Without further improvements in the reconstruction methods, this will result in a large device dependency and poor generalization of the methods.
VII Generalization analysis and effective spectral encoding
We have shown above three important findings in the RGB2HS formulation [7] through extensive experiments. The overfitting issues and metameric failure phenomenon contribute to the generalization difficulties in unseen data, particularly for metamers. The limitations could be mitigated with our proposed metameric augmentation along with aberration modeling as a spectral encoding scheme. Some other spectral encoding methods, such as CASSI, could also benefit from metameric augmentation. In Table IX, we show results on one of the best-performing network MST-L [13] trained the ARAD1K dataset and validated on the other three datasets. All the experiments are carried out with Poisson noise npe = 1000, and aberrations from the double Gauss lens (US20210263286A1). We adopt the CASSI settings in  [13].
As indicated by the results, RGB2HS cannot maintain its performance in PSNR and SAM for data in other datasets, even with metameric augmentation to account for metamers during training (see the RGB2HS columns). With the aid of aberrations and metameric augmentation, the reconstruction performance could be boosted for both standard and metamer data. In addition, evident from the CASSI results without metameric augmentation, CASSI offers overall better reconstruction quality, thanks to its spatial-spectral encoding design rooted in the compressive sensing theory. When metameric augmentation is applied to CASSI, its performance has also been boosted, proving the effectiveness of our metameric augmentation scheme. Note that the reconstruction quality varies among datasets, owing to the different spectral content in each dataset.
Effective spectral encoding is paramount to snapshot spectral imaging. The results of our extensive experiments reiterate the credibility to the computational camera approaches for spectral imaging, including both PSF engineering approaches [36, 9, 15] and compressed sensing methods like CASSI [62, 19]. However, learned reconstruction methods for these approaches also suffer from the same dataset issues as the methods analyzed in this paper, making the collection of large-scale, diverse spectral image data a matter of urgency. Our work reveals the underlying problems shared in most snapshot spectral imaging methods. Once they are well recognized, better solutions can be proposed to overcome these challenges in future works.
VIII Conclusion
@@
Our results systematically demonstrate that it is impossible to accurately reconstruct spectrum solely from RGB images. In order to realize the dream of spectral estimation from arbitrary RGB sources, it is necessary to coherently and jointly diversify the spectral contents in hyperspectral image datasets, adopt side-channel information from the optical system, and embrace versatile spectral data augmentation methods to fully enable the power of networks in adaptation to whole families of aberrations.
@@
-A Image Formation Model
@@
${\mathbf{X}}=\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right],$ (9)
@@
${\mathbf{Q}}=\left[{\mathbf{q}}_{1},{\mathbf{q}}_{2},{\mathbf{q}}_{3}\right]=\left[\begin{matrix}q_{11}&q_{21}&q_{31}\\ q_{21}&q_{22}&q_{32}\\ \vdots&\ddots&\vdots\\ q_{K1}&q_{K2}&q_{K3}\\ \end{matrix}\right],$ (10)
@@
${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$ (11)
@@
${\mathbf{Y}}=\left[{\mathbf{y}}_{1},{\mathbf{y}}_{2},{\mathbf{y}}_{3}\right],$ (12)
@@
${\mathbf{w}}_{k}={\mathbf{A}}_{k}{\mathbf{x}}_{k},\quad k\in[1,2,\dots,K],$ (13)
@@
${\mathbf{W}}$ $=\left[{\mathbf{w}}_{1},{\mathbf{w}}_{2},\dots,{\mathbf{w}}_{K}\right]$ (14)
@@
${\mathbf{A}}=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right],$ (15)
@@
${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right),$ (16)
@@
${\mathbf{A}}{\mathbf{X}}$ $=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right]\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right]$ (17)
@@
${\mathbf{Z}}={\mathbf{W}}{\mathbf{Q}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}}.$ (18)
@@
-B Effect of clipping to non-negative values.